A light rail commuter train decelerates at a rate of 1.65m/s2. If it’s traveling at a max speed of 80km/h, how long does it take to come to a stop? Round two decimal places.


1
Expert's answer
2021-02-15T00:57:53-0500
"t=\dfrac{v-v_0}{a},""t=\dfrac{0-80\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}}{-1.65\ \dfrac{m}{s^2}}=13.47\ s."